# Magnetic Materials for Quantum Magnonics

### v1

@@ II Available magnetic materials
Antiferromagnets have emerged as a promising platform for magnonics and quantum magnonics, driven by a series of recent theoretical advances. Foundational work has shown that antiferromagnetic insulators can support superfluid spin transport, enabling coherent long-range spin currents without net magnetization Takei et al. (2014). Their rich spin textures and ultrafast dynamics further highlight the potential of antiferromagnets for spintronic and magnonic devices Gomonay et al. (2018). More recently, the quantum aspects of antiferromagnetic magnons have been explored. A transmon qubit has been proposed as a probe of the quantum characteristics of antiferromagnetic magnonsAzimi-Mousolou et al. (2023), while spin–orbit interactions have been predicted to mediate their hybridization with plasmons, together opening access to entanglement, nonclassical correlations, and new hybrid excitations with enhanced functionalityDyrdał et al. (2023). A representative antiferromagnet of growing interest in magnonics is hematite ($\alpha$-Fe2O3) and shown in Fig. 1 (a). Bulk studies at GHz frequencies report low Gilbert damping (see Tab. 1) Chen et al. (2025); Hamdi, Posva, and Grundler (2023); Wang et al. (2023b). In the easy-axis phase below the Morin transition, all-electrical spectroscopy reveals nearly gapless antiferromagnetic magnons bound to $180^{\circ}$ domain walls that propagate over micrometre distances with a group velocity of about $6\,\mathrm{km/s}$, as illustrated in Fig. 1 (b). The domain-wall mode exhibits a full width at half maximum (FWHM) frequency linewidth $\Delta f_{\mathrm{FWHM}}\approx 1.2\,\mathrm{GHz}$ and a corresponding lifetime $\tau\approx 0.3\,\mathrm{ns}$ Chen et al. (2025). Another canted antiferromagnet with even better damping characteristics is yttrium orthoferrite ($\mathrm{YFeO_{3}}$) Das et al. (2022). It exhibits linewidths below 10 mT at frequencies around 350 GHz and a temperature of 20 K, corresponding to magnon lifetimes slightly above one nanosecond. Although antiferromagnets can reach terahertz frequencies, which push the quantum regime to higher temperatures (e.g. $1\,\mathrm{THz}\approx 48\,\mathrm{K}$), the relatively short lifetimes currently limit the use of hematite, $\mathrm{YFeO_{3}}$ and other antiferromagnets for quantum magnonics.
@@
Recently, altermagnets have emerged as a distinct class of magnetic materials beyond conventional ferro- and antiferromagnets (see Fig. 1 (c)), characterized by compensated collinear order in real space and non-relativistic alternating spin splitting in momentum space Song et al. (2025b); Gomonay et al. (2024). This unique duality yields terahertz-frequency, field-robust dynamics with negligible stray fields—features reminiscent of antiferromagnets — while also enabling spin-polarized band structures and novel transport phenomena such as the crystal anomalous Hall effect, the spin-splitter effect, and anisotropic magnon spectra. Depending on the orbital symmetry of the momentum-dependent spin splitting, altermagnets are classified into even-parity (d-, g- and i-wave) orders with nodal sign changes (e.g., $d$-wave materials Song et al. (2025b); Zarzuela et al. (2025) like RuO2 Gomonay et al. (2024), MnTeAmin et al. (2024)). Alltogether, this makes altermagnets attractive for quantum magnonics, although strong magnon–magnon interactions can lead to reduced lifetimes, and challenges in thin-film growth together with high damping limits their near-future applicability.
@@
Recently, two-dimensional (2D) van der Waals magnets, illustrated in Fig. 1 (d) and (e), have attracted renewed interest in magnonicsKruglyak et al. (2024). A notable example is the van der Waals antiferromagnet CrPS4 (see Tab. 1 and Fig. 1 (e)), which supports long-distance magnon transportde Wal et al. (2023). Likewise, van der Waals magnets such as CrCl3 exhibit standing spin waves. Experiments resolve modes across a thickness of $20\,\mu\mathrm{m}$, demonstrating control of spin-wave modes in 2D systemsKapoor et al. (2020). Propagating spin waves have also been detected in the 2D van der Waals ferromagnet Fe5GeTe2, despite relatively high magnetic dampingSchulz et al. (2023). These results highlight 2D materials as tunable and versatile platforms for future magnonic applications. Organic magnets also offer promise through chemical tunability and potentially low-damping propagation. For example, V(TCNE)x exhibits low with magnon lifetimes up to 40 nsLiu et al. (2020); McCullian et al. (2020). However, for both van der Waals and organic magnets, the material quality must be improved in order to further increase the lifetime of the magnons required for becoming a medium candidate for single-magnon experiments.
@@
For long magnon lifetimes and extended free paths, it is essential to use electrically insulating materials to eliminate the dissipation channel from magnons into free electrons. A prominent class of such magnetic insulators are the M-typ hexaferrites, with barium hexaferrite (BaFe12O19, BaM) as their flagship representative. BaM is an insulating ferrimagnet with robust performance across a wide temperature range ($T_{\mathrm{C}}\approx 725$ K) Harris (2012), high coercivity and remanence, and a saturation magnetization of $\mu_{0}M_{\mathrm{s}}\approx 380$ kA/m at RT Harris (2012). Its strong intrinsic magnetic anisotropy ($\mu_{0}H_{\mathrm{a}}\approx 1353$ kA/m) enables broadband frequency operation (50–65 GHz in undoped films at low magnetic fields Song et al. (2010), and 20–100 GHz in doped films Levchenko et al. (2024) under moderate biasing magnetic fields <0.8<0.8 T), as well as comparatively low magnetic losses (Gilbert damping constants as low as $\alpha\sim 7\times 10^{-4}$Li et al. (2016); Malkinski (2012). BaM supports SW with high group velocities, while its perpendicular magnetic anisotropy permits isotropic spin-wave propagation in the forward-volume geometry under low magnetic bias. Such advantages make it a highly attractive material for magnonic radio frequency (RF) and data-processing applications (e.g., for 5G FR25G tools for RF Wireless (2025); Levchenko et al. (2024) 24.25 GHz to 71.0 GHz frequency band), yet magnon lifetimes between several and tens of nanoseconds remain insufficient for experiments that require single-magnon operations. Furthermore, BaM is still mostly limited to macro- and microscale applications due to the lack of a reliable methodology for reproducible growth of high-quality nm-thick films, although ongoing efforts aim to overcome this challenge.
@@
The longest magnon lifetimes have so far been achieved in YIG, which remains the only material with sufficiently low damping to sustain magnons long enough for single-magnon excitation, measurement, and manipulation. This outstanding material is a ferrimagnetic insulatorCherepanov, Kolokolov, and L’vov (1993); Serga, Chumak, and Hillebrands (2010) has a well understood magnetization behaviorSerha et al. (2025a) and magnetic anisotropyBobkov and Zavislyak (1997). The damping of coherent magnons arises primarily from scattering on crystallographic and surface defectsSpencer, LeCraw, and Clogston (1959); Melkov et al. (2004), from rare-earth impuritiesMihalceanu et al. (2018); Dillon and Nielsen (1959); Spencer, LeCraw, and Linares (1961); Serha et al. (2025b), and from interactions with thermal magnonsPincus, Sparks, and LeCraw (1961); White and Sparks (1963) (spin-spin interaction) and phononsKasuya and LeCraw (1961); Rückriegel et al. (2014); Streib et al. (2019) (spin-lattice interaction). As summarized in Tab. 2, for example at millikelvin temperatures an FMR linewidth of $0.02\,\mathrm{mT}$ corresponds to a magnon lifetime of about $300\,\mathrm{ns}$Serha et al. (2025b), i.e., the minimally required timescale of a few hundred nanoseconds for single-magnon experimentsLachance-Quirion et al. (2020); Xu et al. (2023). Even so, lifetimes of this order are still sufficient to sustain spin-wave coherence for more than 10 µs Makiuchi et al. (2024) and to support magnon propagation over distances exceeding 100 µm using a 110 nm-thin YIG LPE film.Bensmann et al. (2025).
@@
For advanced quantum magnonic operations, longer magnon lifetimes are required. Previous experiments used the Kittel mode, a uniform precession of the magnetization across the whole sample. Its damping is dominated by two-magnon scattering from lattice and surface defects and is therefore largely temperature independent. On cooling to millikelvin temperatures, the lifetime decreases only slightly compared with RTKosen et al. (2019); Serha et al. (2025b). In contrast, short-wavelength magnons with finite wave vector that are localized inside the sample are insensitive to surface quality. For these dipolar-exchange spin waves (DESW) the dominant dissipation is scattering from thermal magnons and phonons. Recently, Serha et al.Serha et al. (2025b) showed that when this thermal bath is depleted the lifetime of short-wavelength magnons increases by more than an order of magnitude and reaches up to 18 µs. These magnons were generated through the intrinsic nonlinearity of the system via three-magnon splitting, with the lifetime set by the power threshold for entering the nonlinear regimeSuhl (1957).
@@
The lifetime of these secondary magnons increases drastically as the temperature decreases from a few kelvin to the millikelvin range as shown in Fig. 1 (f). Multimagnon scattering on thermal magnons and interactions with thermal phonons vanish as the temperature approaches absolute zero. Surface and lattice defects that give rise to two-magnon scattering do not significantly affect DESW, since these bulk modes have micrometer-scale wavelengths. In theory, the lifetime of these magnons should diverge at millikelvin temperatures. In practice, however, it saturates below about 250 mKSerha et al. (2025b). Interestingly, this saturation occurs at different lifetime levels, with crystal contamination with rare and heavy metal ions becoming the limiting factor. Higher-purity samples exhibit longer lifetimes, indicating that values beyond 20 µs — the dephasing time $T_{2}$ of state-of-the-art superconducting qubits Burnett et al. (2019) — are within reach when YIG is grown from specially purified yttrium oxide ($\mathrm{Y_{2}O_{3}}$), which contains most of the rare-earth and heavy-earth impurities responsible for limiting YIG quality at low temperatures.
@@ III YIG films for quantum magnonics
Property bulk YIGSerha et al. (2025b); Maier-Flaig et al. (2017); Klingler et al. (2017) YIG/GGGSerha et al. (2025d) YIG/YSGGYoussef et al. (2025) YIG/YSGAGSerha et al. (2025d)
@@
Gilbert damping $\alpha$ @ RT $3\times 10^{-5}$ $4.3\times 10^{-5}$ $1.0\times 10^{-4}$ $4.3\times 10^{-5}$
@@
Figure 2: FMR FWHM linewidth $\Delta B$ (measure for magnetic damping) as a function of temperature $T$ for an FMR frequency $f_{\mathrm{FMR}}\approx 8\,\mathrm{GHz}$, displayed on a double logarithmic scale. The graph compares measurements obtained in this work with selected results of sputtered and LPE grown YIG films from relevant literature, enabling a direct evaluation of temperature-dependent damping behavior across different samples, their substrates and studies Will-Cole et al. (2023); Guo et al. (2023); Legrand et al. (2025); Serha et al. (2025d) (taken and adapted from Serha et al. (2025d)).
@@
To use magnons as quantum data carriers in quantum magnonics, it is crucial to move beyond standing magnons in spherical resonatorsLachance-Quirion et al. (2019, 2020); Xu et al. (2023); Li et al. (2022); Rani et al. (2025) and establish YIG films as the active medium for propagating single magnons. At RT, the highest bulk like quality YIG films, from micrometer to nanometer scales, have been grown on paramagnetic GGG substrates because of the nearly lattice matching between the two crystals Dubs et al. (2017, 2020); Ding et al. (2020); Heyroth et al. (2019); Cornelissen et al. (2015); Onbasli et al. (2014); Hahn et al. (2014); Askarzadeh and Shokrollahi (2025).
@@
However, quantum magnonics must operate at very low temperatures down to the millikelvin range. Under these conditions, paramagnetic substrates such as GGG become readily magnetized by an externally applied magnetic field Danilov et al. (1989); Serha et al. (2024). The magnetized substrate generates an inhomogeneous stray field that depends on its geometry and imposes a strong gradient in the internal field of the YIG film. For example, a common $(5\times 5\times 0.5)$ mm3 GGG substrate at 1.8 K and an in-plane bias field of 600 mT produces stray fields of up to 60 mT at the edges and about 12 mT at the center opposing the external fieldSerha et al. (2024). This gradient is already broadening the FMR linewidths measured over the sampleSerha et al. (2025c). By selecting an appropriate substrate geometry or by microstructuring the YIG film, the GGG-induced stray-field gradient can be eliminated across the magnetic medium, resulting in a homogeneous internal field in YIGSchmoll et al. (2025b). However, the partially magnetically ordered substrate couples to the YIG spin system Schmoll et al. (2025a), opening extra dissipation channels that broaden the ferromagnetic resonance (FMR) linewidth Serha et al. (2025c, 2024); Guo et al. (2022); Jermain et al. (2017); Mihalceanu et al. (2018); Kosen et al. (2019) and reduce magnon lifetimes of propagating dipolar magnons Schmoll et al. (2025a). The linewidth broadening effect is shown in Fig. 1 (g)Serha et al. (2025c) and Tab. 2 reports a linewidth difference of more than a factor of four between RT and $T\,\rightarrow 0$, even in the absence of the inhomogeneous GGG-induced stray field. Such a decrease in magnon lifetime undermines the advantages that YIG would otherwise offer for quantum-magnonic experiments. MicrostructuringSchmoll et al. (2025b) and operating at low external fieldsKnauer et al. (2023) can mitigate the stray-field gradient to some extent. Nevertheless, eliminating the coupling between the substrate spin system and the YIG film requires the use of diamagnetic substrates. For a comparison of the temperature-dependent damping behavior between YIG films grown on GGG and on diamagnetic alternative substrates, Fig. 2 presents literature results on sputtered and LPE-grown YIG films from different studies, covering various samples and substratesWill-Cole et al. (2023); Guo et al. (2023); Legrand et al. (2025); Serha et al. (2025d).
@@
Yttrium aluminum garnet (YAG) was among the first diamagnetic substrates proposed to replace paramagnetic GGG in YIG heterostructures, aiming to unlock new possibilities for quantum magnonics at cryogenic temperatures by avoiding substrate magnetization in applied fields. However, despite the drawbacks of GGG at low temperatures, YIG/GGG still outperforms YIG/YAG at both RT and low temperature (see Fig. 2), primarily because the large lattice mismatch between YIG and YAG degrades crystalline quality of the YIG filmSposito et al. (2013); Krysztofik et al. (2021); Serha et al. (2025d).

### v2

@@ II Available magnetic materials
Antiferromagnets have emerged as a promising platform for magnonics and quantum magnonics, driven by a series of recent theoretical advances. Foundational work has shown that antiferromagnetic insulators can support superfluid spin transport, enabling coherent long-range spin currents without net magnetization Takei et al. (2014). Their rich spin textures and ultrafast dynamics further highlight the potential of antiferromagnets for spintronic and magnonic devices Gomonay et al. (2018). More recently, the quantum aspects of antiferromagnetic magnons have been explored. In this context, a transmon qubit has been proposed as a probe of their quantum characteristicsAzimi-Mousolou et al. (2023). In parallel, spin–orbit interaction has been predicted to mediate the hybridization of antiferromagnetic magnons with plasmons, i.e., collective charge-density oscillations of conduction electrons, resulting in coherently coupled magnon–plasmon modes with modified dispersions and nontrivial band topology described within a quantum-mechanical frameworkDyrdał et al. (2023).
A representative antiferromagnet of growing interest in magnonics is hematiteGückelhorn et al. (2022); El Kanj et al. (2023); Gückelhorn et al. (2023) ($\alpha$-Fe2O3) and shown in Fig. 1 (a). Bulk studies at GHz frequencies report low Gilbert damping (see Tab. 1) Chen et al. (2025); Hamdi et al. (2023); Wang et al. (2023a). In the easy-axis phase below the Morin transition, all-electrical spectroscopy reveals nearly gapless antiferromagnetic magnons bound to $180^{\circ}$ domain walls that propagate over micrometre distances with a group velocity of about $6\,\mathrm{km/s}$, as illustrated in Fig. 1 (b). The domain-wall mode exhibits a full width at half maximum (FWHM) frequency linewidth $\Delta f_{\mathrm{FWHM}}\approx 1.2\,\mathrm{GHz}$ and a corresponding lifetime $\tau\approx 0.3\,\mathrm{ns}$ Chen et al. (2025). Another canted antiferromagnet with even better damping characteristics is yttrium orthoferrite ($\mathrm{YFeO_{3}}$) Das et al. (2022). It exhibits linewidths below 10 mT at frequencies around 350 GHz and a temperature of 20 K, corresponding to magnon lifetimes slightly above one nanosecond. Although antiferromagnets can reach terahertz frequencies, which push the quantum regime to higher temperatures (e.g. $1\,\mathrm{THz}\approx 48\,\mathrm{K}$), the relatively short lifetimes currently limit the use of hematite, $\mathrm{YFeO_{3}}$ and other antiferromagnets for quantum magnonics.
@@
Recently, altermagnets have emerged as a distinct class of magnetic materials beyond conventional ferro- and antiferromagnets (see Fig. 1 (c)), characterized by compensated collinear order in real space and non-relativistic alternating spin splitting in momentum space Song et al. (2025a); Gomonay et al. (2024). This unique duality yields terahertz-frequency, field-robust dynamics with negligible stray fields—features reminiscent of antiferromagnets — while also enabling spin-polarized band structures and novel transport phenomena, such as the crystal anomalous Hall effect, the spin-splitter effect, and anisotropic magnon spectra. Depending on the orbital symmetry of the momentum-dependent spin splitting, altermagnets are classified into even-parity (d-, g- and i-wave) orders with nodal sign changes (e.g., d-wave materials Song et al. (2025a); Zarzuela et al. (2025) like RuO2 Gomonay et al. (2024), MnTeAmin et al. (2024)). They are distinguished from s‑ and p‑wave (odd or no‑node) patterns found in ferromagnets or antiferromagnets. A prominent example is hematite ($\alpha$-Fe2O3), a well-established antiferromagnet in magnonicsChen et al. (2025); Hamdi et al. (2023); Wang et al. (2023a); El Kanj et al. (2023); Gückelhorn et al. (2023), which has recently been identified as an altermagnet and shown to host anisotropic magnon excitations consistent with momentum-dependent spin splittingHoyer et al. (2025); Galindez‐Ruales et al. (2025). Altogether, this makes altermagnets attractive for quantum magnonics, although the reduced lifetimes due to magnon–magnon interactionsGarcia-Gaitan et al. (2025) should be taken into account, as is the case for antiferromagnets.
@@
[…]
Organic magnets also offer promise through chemical tunability and potentially low-damping propagation. For example, V(TCNE)x exhibits lifetimes up to 40 nsLiu et al. (2020); McCullian et al. (2020). Low magnetic damping, together with their other advantageous properties, makes these materials highly promising for quantum magnonics. Further research is needed to fully understand and explore their potential.
@@
For long magnon lifetimes and extended free paths, it is essential to use electrically insulating materials to eliminate the dissipation channel from magnons into free electrons. A prominent class of such magnetic insulators are the M-type hexaferrites, with barium hexaferrite (BaFe12O19, BaM) as their flagship representative. BaM is an insulating ferrimagnet with robust performance across a wide temperature range ($T_{\mathrm{C}}\approx 725$ K) Harris (2012), high coercivity and remanence, and a saturation magnetization of $M_{\mathrm{s}}\approx 380$ kA/m at RT Harris (2012). Its strong intrinsic magnetic anisotropy ($H_{\mathrm{a}}\approx 1353$ kA/m) enables broadband frequency operation (50–65 GHz in undoped films at low magnetic fields Song et al. (2010), and 20–100 GHz in doped films Levchenko et al. (2026) under moderate biasing magnetic fields <0.8<0.8 T), as well as comparatively low magnetic losses (Gilbert damping constants as low as $\alpha\sim 7\times 10^{-4}$Li et al. (2016); Malkinski (2012).
BaM supports spin waves with high group velocities, while its strong perpendicular magnetic anisotropy permits isotropic forward-volume spin-wave propagation in the film plane already under low external magnetic bias. Combined with intrinsically low magnetic damping, with Gilbert damping constants on the order of $\alpha\sim 10^{-4}$, these properties make BaM a highly attractive material for magnonic radio-frequency (RF) signal processing and data-processing applications, including operation in the 5G FR2 frequency range (24.25–71.0 GHz) 5G tools for RF Wireless (2025); Levchenko et al. (2026). BaM uniquely enables magnonic access to technologically relevant frequency bands above 60 GHz that are difficult to reach with other ferrimagnets. These include the 60 GHz band used for in-cabin automotive radar and health-monitoring applications Infineon60GHz (2024), as well as the 76–81 GHz band employed in long-range automotive radar systems IB-Lenhardt (2024). Importantly, the large uniaxial anisotropy enables self-magnetized operation, allowing access to high-frequency magnonic modes with little or no external bias field. In nm-thick BaM waveguides, the fundamental spin-wave frequencies naturally lie in the tens-of-GHz range. These frequencies are primarily set by the internal anisotropy field rather than an applied magnetic field.
By providing high-frequency and potentially field-free spin-wave operation, BaM positions magnonics as a viable and competitive platform for next-generation RF signal processing and sensing technologies, enabling contributions to application domains that remain inaccessible to conventional magnonic materials. At the same time, the intrinsically high operating frequencies and low damping of BaM make this material class a promising candidate for high-frequency quantum magnonics, where large magnon energies are advantageous for suppressing thermal occupation in quantum magnonics. The main limitation of BaM for modern magnonics arises from the absence of lattice-matched growth techniques capable of producing high-quality single-crystal nanoscale films that simultaneously provide low magnetic loss, low bias fields, and high-frequency performance. Systematic studies of BaM at cryogenic temperatures are therefore highly desirable and remain largely unexplored.
Europium Chalcogenides
A class of magnetic insulators and semiconductors that become magnetically ordered only at cryogenic temperatures — and are therefore unsuitable for room-temperature magnonics and spintronics, yet potentially relevant for quantum magnonics — are the europium chalcogenides (Eu2+X2-)Wachter (1979); Schlipf et al. (2013); Boncher et al. (2015); Dietl et al. (2019). This material family includes europium oxide (EuO), europium sulfide (EuS), europium selenide (EuSe), and europium telluride (EuTe). All compounds crystallize in the rock-salt structure and exhibit band gaps ranging from approximately 1.2 eV (EuSe) to about 2.4 eV (EuTe), making them model systems for studying localized Heisenberg magnetism driven by 4f electronsKolokolov et al. (2025).
The magnetic order in europium chalcogenides is governed by nearest- and next-nearest-neighbor exchange interactions, resulting in ferromagnetic ground states in EuO and EuS with Curie temperatures of $T_{\mathrm{C}}\approx 69\,\mathrm{K}$ and $T_{\mathrm{C}}\approx 16.6\,\mathrm{K}$, respectively. In contrast, EuSe exhibits a complex antiferromagnetic ground state below $T_{\mathrm{N}}\approx 4.6\,\mathrm{K}$ with field-induced metamagnetic phases that depend sensitively on temperature and external magnetic field, while EuTe is an antiferromagnet with a Néel temperature of $T_{\mathrm{N}}\approx 9.6\,\mathrm{K}$. Owing to the large magnetic moment of the Eu2+ ion ($S=7/2$), these materials exhibit exceptionally high saturation magnetizations, reaching approximately 1900 kA/m in EuO and about 1300 kA/m in EuTe (in the fully field-saturated antiferromagnetic state). As a result, spin-wave frequencies in europium chalcogenides are significantly higher than in most conventional magnetic insulators.
Magnetic damping and resonance linewidths in europium chalcogenides have been investigated primarily at low temperatures using ferromagnetic and antiferromagnetic resonance techniques. In high-quality single crystals, remarkably narrow linewidths have been reported, with values on the order of millitesla for EuS, while comparable or slightly larger low-temperature linewidths have been observed in EuO, EuSe, and EuTe depending on the magnetic phase and field configurationvon Molnár and Lawson (1965); Franzblau et al. (1967); Brown et al. (1968); Battles and Everett (1969); Eastman (1968); Kunii et al. (1974); Dillon and Olsen (1964). Taking the lowest reported linewidth for EuS, $\Delta B\approx 1\,\mathrm{mT}$ at $T\approx 1.4\,\mathrm{K}$ and microwave frequencies around 22 GHz, yields a magnon lifetime of approximately $\tau\approx 6\,\mathrm{ns}$Franzblau et al. (1967).
EuS has recently been employed in studies of magnon transport driven by the spin Seebeck effect in ferromagnetic EuS thin filmsAguilar-Pujol et al. (2023), demonstrating its relevance for cryogenic magnonics. While the magnon lifetimes in europium chalcogenides are currently much shorter than those achieved in quantum magnonics experiments based on YIG, they are already comparable to coherence times relevant for hybrid quantum systems operating in the few-gigahertz regime. Given the localized 4f magnetism, the absence of itinerant charge carriers, and the presence of strong exchange fields, further improvements in linewidth and magnon lifetime may be expected in defect-reduced samples and at millikelvin temperatures. These considerations position europium chalcogenides as a promising, yet largely unexplored, materials platform for quantum magnonics and motivate systematic investigations of their magnon lifetimes, coherence properties, and dissipation mechanisms in the quantum limit.
@@
The longest magnon lifetimes have so far been achieved in YIG, which remains the only material particularly used for single-magnon excitation, measurement, and manipulation. This outstanding material is a ferrimagnetic insulatorCherepanov et al. (1993); Serga et al. (2010) has a well understood magnetization behaviorSerha et al. (2025c) and magnetic anisotropyBobkov and Zavislyak (1997). The damping of coherent magnons arises primarily from scattering on crystallographic and surface defectsSpencer et al. (1959); Melkov et al. (2004), from rare-earth impuritiesSerha et al. (2025b); Mihalceanu et al. (2018); Pardavi-Horvath (2000); Dillon and Nielsen (1959); Spencer et al. (1961), and from interactions with thermal magnonsPincus et al. (1961); White and Sparks (1963) (spin-spin interaction) and phononsKasuya and LeCraw (1961); Klein et al. (2004); Naletov et al. (2007); Rückriegel et al. (2014); Streib et al. (2019); Müller et al. (2024); Ba et al. (2025); Cherkasskii et al. (2025) (spin-lattice interaction). As summarized in Tab. 2, for example at millikelvin temperatures, an FMR linewidth of $0.02\,\mathrm{mT}$ corresponds to a magnon lifetime of about $300\,\mathrm{ns}$Serha et al. (2025b), i.e., the minimally required timescale of a few hundred nanoseconds for single-magnon experimentsLachance-Quirion et al. (2020); Xu et al. (2023). Even so, lifetimes of this order are still sufficient to sustain spin-wave coherence for more than 10 µs Makiuchi et al. (2024) and to support magnon propagation over distances exceeding 100 µm using a 110 nm-thin YIG liquid-phase epitaxy (LPE) film.Bensmann et al. (2025).
@@
For advanced quantum magnonic operations, longer magnon lifetimes are required. Previous experiments used the Kittel mode, a uniform precession of the magnetization across the whole sample. The damping of the FMR mode and dipolar magnons with large wavelengths is dominated by two-magnon scattering from lattice and surface defects, and is therefore largely temperature independent. On cooling to millikelvin temperatures, the lifetime decreases only slightly compared with RTKosen et al. (2019); Schmoll et al. (2025b); Serha et al. (2025b). In contrast, short-wavelength magnons with finite wave vector, that are localized inside the sample, are insensitive to surface quality. For these dipolar-exchange spin waves (DESW) the dominant dissipation is scattering from thermal magnons and phonons. Recently, Serha et al.Serha et al. (2025b) showed that when this thermal bath is depleted the lifetime of short-wavelength magnons increases by more than an order of magnitude and reaches up to 18 µs. These magnons were generated through the intrinsic nonlinearity of the system via three-magnon splitting, with the lifetime set by the power threshold for entering the nonlinear regimeSuhl (1957).
The lifetime of these secondary magnons increases drastically as the temperature decreases from a few kelvin to the millikelvin range, as shown in Fig. 1 (f). As the temperature approaches absolute zero, multimagnon scattering on thermal magnons and interactions with thermal phonons vanish, and because short-wavelength DESW remain localized within the crystal volume and are effectively decoupled from dipolar coupling to surrounding dissipation channels, their intrinsic relaxation is strongly suppressed, resulting in a pronounced enhancement of the magnon lifetime at millikelvin temperaturesSerha et al. (2025b).
@@
In theory, the lifetime of these magnons is expected to approach infinity at millikelvin temperatures. In practice, however, it saturates below about 250 mKSerha et al. (2025b). Interestingly, this saturation occurs at different lifetime levels, with crystal contamination with rare and heavy metal ions becoming the limiting factor. This impurity-induced lifetime-limiting bottleneck is evident in Fig. 1 (f), where a reduction in impurity concentration from Sphere 1 to Sphere 3 leads to an increase in the saturation lifetime of DESW magnons from approximately 5 µs to 18 µs. As higher-purity samples exhibit longer magnon lifetimes, indicating that values beyond 20 µs — the dephasing time $T_{2}$ of state-of-the-art superconducting qubits Burnett et al. (2019) — are within reach when YIG is grown from specially purified yttrium oxide ($\mathrm{Y_{2}O_{3}}$), which contains most of the rare-earth and $L$-state transition metal ion responsible for limiting YIG quality at low temperatures.
@@ III YIG films for quantum magnonics
Property bulk YIGSerha et al. (2025b); Maier-Flaig et al. (2017); Klingler et al. (2017) YIG/GGGSerha et al. (2025a) YIG/YSGGYoussef et al. (2025) YIG/YSGAGSerha et al. (2025a)
@@
Gilbert damping $\alpha$ @ RT ($\times 10^{-5}$) $3\pm 0.5$ $4.3\pm 0.4$ $10\pm 1$ $4.3\pm 0.6$
@@
Figure 2: FMR FWHM linewidth $\Delta B$ (measure for magnetic damping) as a function of temperature $T$ for an FMR frequency $f_{\mathrm{FMR}}\approx 8\,\mathrm{GHz}$, displayed on a double logarithmic scale. The graph compares measurements obtained in this work with selected results of sputtered (hollow symbols) and LPE grown (filled symbols) YIG films from relevant literature, enabling a direct evaluation of temperature-dependent damping behavior across different samples, their substrates and studies Will-Cole et al. (2023); Guo et al. (2023); Legrand et al. (2025); Serha et al. (2025a). Reproduced from R. O. Serha et al., arXiv:2508.19044 (2025); licensed under a Creative Commons Attribution-NonCommercial-NoDerivatives 4.0 International ($C$ BY-NC-ND) license.
@@
To use magnons as quantum data carriers in quantum magnonics, it is crucial to move beyond standing magnons in spherical resonatorsLachance-Quirion et al. (2019, 2020); Xu et al. (2023); Li et al. (2022); Rani et al. (2025) and establish YIG films as the active medium for propagating single magnons. At RT, the highest bulk-like quality YIG films, from micrometer to nanometer scales, have been grown on paramagnetic GGG substrates because of the nearly perfect lattice matching between the two crystals Dubs et al. (2017, 2020); Ding et al. (2020); Heyroth et al. (2019); Cornelissen et al. (2015); Onbasli et al. (2014); Hahn et al. (2014); Askarzadeh and Shokrollahi (2025).
@@
However, quantum magnonics must operate at very low temperatures down to the millikelvin range. Under these conditions, paramagnetic substrates, such as GGG become readily magnetized by an externally applied magnetic field Danilov et al. (1989); Serha et al. (2024). Counterintuitively, the GGG substrate does not behave as a simple paramagnet describable by a Brillouin function at temperatures below 500 mK. Instead, in this temperature range it exhibits a complex magnetic phase diagram and enters a frustrated magnetic state, in which the magnetization becomes nearly temperature independent and is governed primarily by the external magnetic fieldSerha et al. (2024); Deen et al. (2015); Petrenko et al. (1998); Schiffer et al. (1994). The magnetized substrate generates an inhomogeneous stray field that depends on its geometry and imposes a strong gradient in the internal field of the YIG film. For example, a common $(5\times 5\times 0.5)$ mm3 GGG substrate at 1.8 K and an in-plane bias field of 600 mT produces stray fields of up to 60 mT at the edges and about 12 mT at the center opposing the external fieldSerha et al. (2024). This gradient is already broadening the FMR linewidths measured over the sampleSerha et al. (2025d). By selecting an appropriate substrate geometry or by microstructuring the YIG film, the GGG-induced stray-field gradient can be eliminated across the magnetic medium, resulting in a homogeneous internal field in YIGSchmoll et al. (2025a). However, the partially magnetically ordered substrate couples to the YIG spin system Schmoll et al. (2025b), opening extra dissipation channels that broaden the ferromagnetic resonance (FMR) linewidth Serha et al. (2025d, 2024); Guo et al. (2022); Jermain et al. (2017); Mihalceanu et al. (2018); Kosen et al. (2019) and reduce magnon lifetimes of propagating dipolar magnons Schmoll et al. (2025b). The linewidth broadening effect is shown in Fig. 1 (g)Serha et al. (2025d) and Tab. 2 reports a linewidth difference of more than a factor of four between RT and $T\,\rightarrow 0$, even in the absence of the inhomogeneous GGG-induced stray field. Such a decrease in magnon lifetime undermines the advantages that YIG would otherwise offer for quantum-magnonic experiments. MicrostructuringSchmoll et al. (2025a) and operating at low external fieldsKnauer et al. (2023) can mitigate the stray-field gradient to some extent. Nevertheless, eliminating the coupling between the substrate spin system and the YIG film requires the use of diamagnetic substrates. For a comparison of the temperature-dependent damping behavior between YIG films grown on GGG and on diamagnetic alternative substrates, Fig. 2 presents literature results on sputtered and LPE-grown YIG films from different studiesWill-Cole et al. (2023); Guo et al. (2023); Legrand et al. (2025); Serha et al. (2025a).
@@
Yttrium aluminum garnet (YAG) was among the first diamagnetic substrates proposed to replace paramagnetic GGG in YIG heterostructures, aiming to unlock new possibilities for quantum magnonics at cryogenic temperatures by avoiding substrate magnetization in applied fields. However, despite the drawbacks of GGG at low temperatures, YIG/GGG still outperforms YIG/YAG at both RT and low temperature (see Fig. 2), primarily because the large lattice mismatch between YIG and YAG degrades crystalline quality of the YIG filmSposito et al. (2013); Krysztofik et al. (2021); Serha et al. (2025a).